OpenStudy (sneville):

What is the solution of the system? Use substitution. 3x + 2y = 7 y = –3x + 11 A. (6, –3) B. (6, –7) C. D. (5, –4)

OpenStudy (dopeeaf):

3x+2y=7 -------------(1) y= -3x + 11 ----------(2) subs (2) into (1) 3x + 2( -3x +11) = 7 3x + (-6x + 22) = 7 3x - 6x + 22 = 7 -3x = -15 x = 5 subs x = 5 into (2) y = -3(5) + 11 y = -15 +11 y = -4 x = 5, y = -4

OpenStudy (questionerpersoner):

Think about it if u put if it was (A) 3x6=18 and 2x-3=-6 18+-6 would equal 12